The 25kg box is stationary on a horizontal surface
lies. Push 75N from the side to move the box
required. Once you move out of the place, continue regular
you need to constantly push it at 60N to reach the top speed. Boxes and
static and kinetic friction between the horizontal surfaces
Find the coefficients.
1
Expert's answer
2020-03-25T11:16:37-0400
"\mu_s=\frac{75}{25\cdot 9.8}=0.306"

"\mu_k=\frac{60}{25\cdot 9.8}=0.245"